OpenStudy (my4babiesandme):

simplify by removing factors of 1. 7m^2+7m/63 m^2+49m

jimthompson5910 (jim_thompson5910):

\[\Large \frac{7m^2+7m}{63m^2+49m}\] \[\Large \frac{7m(m+1)}{63m^2+49m}\] \[\Large \frac{7m(m+1)}{7m(9m+7)}\] \[\Large \frac{\cancel{7m}(m+1)}{\cancel{7m}(9m+7)}\] \[\Large \frac{m+1}{9m+7}\] So \[\Large \frac{7m^2+7m}{63m^2+49m}=\frac{m+1}{9m+7}\]